3(2x + 11) = 9 Reason for each step

6x + 33 = 9 Reason for each step

6x=-24 reason for each step

X=-4 reason for each step


44 - 2(3x + 4) = -18x Reason for each step


44 - 6x - 8 = -18x Reason for each step

-6x + 36 = -18x Reason for each step

36 = -12x Reason for each step


-3 = x Reason for each step
1
Expert's answer
2020-09-25T11:04:49-0400

"ONE\\1.\; 3(2x + 11) = 9; first\ we \;open\ the\ brackets\\ to\ do\ this\ we\ multiply\ each\ term\\ in\ the\ brackets\ by\ a \;common \;multiplier\\2.\;6x + 33 = 9 ;\;each\ term\ enclosed\ in\\ parentheses\\ must\ be\ multiplied\ by\ a\ common\ factor.\\3.\;6x=-24;\;move\ the \;known\ term\ to\ the\\ right\ by\ changing \ the \;plus\ sign\ to\ minus.\\4.\;x=-4; divide\ the\ right\ and\\ left\ sides\ of\ the\ equation\ by\;6\\the \;solution\\ of\ the \;equation\;is\ finished\\TWO\\44 - 2(3x + 4) = -18x;\ we \;open\ the\ brackets\\ to\ do\ this\ we\ multiply\ each\ term\\ in\ the\ brackets\ by\ a \;common \;multiplier\\
44 - 6x - 8 = -18x;now\ we\ subtract\ 8 from\ 44 \\ 
-6x + 36 = -18x \\and \;move\ the\ unknown\\ quantity\ to\ the\ right\\ while\ subtracting\\
36 = -12x\\ divide\ the\ right\ and\\ left\ sides\ of\ the\ equation\ by\;12\\
-3 = x;changing \;the\ right\ part\ to\ the\ left \\x=-3 ; that\ is\ all"